# Domain decomposition-based coupling of Operator Inference Reduced Order Models via the Schwarz alternating method

### v2

@@ 2.2.2 Operator Inference
Operator inference is a form of projection-based model that uses a POD basis to reduce the dimension of the operators assembled for the FOM, which, in this paper, is the heat equation described in (6). In a classical Galerkin ROM, the operator $\mathbf{K}_{r}\in\mathbb{R}^{r\times r}$ is the operator which best represents the action of $\mathbf{K}$ in the reduced space of dimension $r$ defined by the basis $\mathbf{\Psi}_{r}$. That is,
$\mathbf{K}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{K}\mathbf{\Psi}_{r}.$
This is intrusive because it requires access to the FOM matrix $\mathbf{K}$, and, practically, the code that produced $\mathbf{K}$. Similarly, we can obtain $\mathbf{x}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{x}$ and $\mathbf{f}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{f}$, both belonging to $\mathbb{R}^{r}$.
@@
For the problem considered herein, the 2D heat equation, it is straightforward to see that a projection-based ROM will preserve the linear algebraic structure of the FOM, (6). In order to obtain the reduced operators in (10), one solves a regression problem for these reduced operators, $\hat{\mathbf{K}}\in\mathbb{R}^{r\times r}$ and $\hat{\mathbf{B}}\in\mathbb{R}^{r\times m}$. In particular, given $j<\tau$ steps of FOM training data $\mathbf{x}_{i}$, $i=1,\dots,j$, minimize,
$\min_{\hat{\mathbf{K}},\hat{\mathbf{B}}}\sum_{i=1}^{j}\|\dot{\mathbf{x}}_{r,i}-\hat{\mathbf{K}}\mathbf{x}_{r,i}-\hat{\mathbf{B}}\mathbf{g}_{i}-\mathbf{f}_{r,i}\|_{2}^{2}.$ (9)
The subscript $r$ is here to note the components in the ROM dimension which are explicitly known within the training data. In the case that a technique known as re-projection is used, it can be shown that the learned OpInf ROM system recovers the standard Galerkin projected intrusive ROM [25]. If $\dot{\mathbf{x}}_{r,i}$ is not explicitly available, it may be estimated from available data using a difference method. The hat notation is reserved for those components which belong to the OpInf method. As this is a data-driven minimization problem, it is not strictly necessary to use the actual known boundary condition $\mathbf{g}_{i}$ or forcing function $\mathbf{f}_{r,i}$; some approximation or ansatz can be learned instead, and/or these terms can be combined into a single input term.
Once problem (9) is solved, which in implementation will typically involve a QR based approach. The ROM ODE analogue of (6) takes the form
$\dot{\hat{\mathbf{x}}}=\hat{\mathbf{K}}\hat{\mathbf{x}}+\hat{\mathbf{B}}\mathbf{g}+\mathbf{f}_{r},$ (10)
where we now solve for the reduced state variable $\hat{\mathbf{x}}\in\mathbb{R}^{r}$ because the outputs of this ODE are not guaranteed to be the same as those for the standard intrusive Galerkin projection. It is guaranteed that, under the assumptions that the time-stepping scheme is convergent as $dt\to 0$, the approximation $\dot{\hat{\mathbf{x}}}_{j}$ converges to the true time derivative at time $t_{j}$ as $dt\to 0$, and linear independence of supplied data ($d\geq\tau$, the total number of snapshots), then, for all $0<\epsilon\in\mathbb{R}$ there exists a timestep $dt$ for the FOM and a $r\leq d$ such that,
@@
The same result holds for $\hat{\mathbf{B}}$ [26].
While this overview has focused on the heat equation, we again stress that operator inference is general to low-order polynomial non-linearity, with results previously shown for a polynomial non-linearity of degree 3 in a one-dimensional (1D) nuclear reactor model in [26], for a 2D single injector combustion model in [36], and for a three-dimensional (3D) rotating detonation rocket engine in [9].
@@ 3 Software Implementations
The Schwarz coupling implementations we will test are FOM-ROM and ROM-ROM, where FOM refers to a finite element (FE) simulation and ROM to operator inference. We will draw our data source from a monolithic FE simulation, but the FOM-ROM coupling will have its simulation run from scratch on its own subdomain independently of the monolithic simulation.
@@
We use the open source FE library FEniCSx in Python for all FE simulations. FEniCSx is a FE problem solving environment with the ability to define variational forms close to actual mathematical notation [1], along with various tools [2, 35, 34] to enable the creation of a FE space on which computations can be performed. Boundary conditions are enforced strongly through a lifting method leading to a FOM system similar to (6).
@@
Operator inference is performed with the Python library OpInf, implemented by Shane McQuarrie [21]. OpInf is a set of tools to carry out POD and infer the required operators as described in Section 2.2. The required data will be drawn from a monolithic FE simulation on $\Omega$, also performed in FEniCSx. When solving the OpInf regression problem (9), we estimate the required state derivative information $\dot{\hat{\mathbf{x}}}$ for through a first order backward difference of available state data, and our ROM ODE (10) will be solved using backward Euler.
@@
The numerical stability of OpInf models is a significant concern and area of active research. Operator inference models are not guaranteed to be structure-preserving [11], and one key structure that we care about is the theoretical stability of the ODE system (10). FOM solutions are generated presently through the FE method, and as such the eigenvalues of $\mathbf{K}$ in (6) are guaranteed to be negative. This structure is not guaranteed to be passed on to $\hat{\mathbf{K}}$, and often is not. There are a variety of strategies that can help improve stability of reduced models – independently of OpInf, data pre-processing is commonly used in ROMs to create a model for a transformed set of snapshots. We will typically use a centering approach where we build a model for snapshot data $\mathbf{x}-\mathbf{\bar{x}}$, the mean-centered snapshots.
@@
Specific to OpInf, we also employ a regression regularization strategy for problem (9). In literature [26], the basic OpInf problem is typically cast into the form,
$\min_{\mathbf{O}}\|\mathbf{D}\mathbf{O}^{T}-\mathbf{R}^{T}\|_{F},$ (11)
[…]
@@
Finally, the Schwarz method that couples these problems has been implemented by the authors of this paper in accordance with Section 2.1. To simplify the boundary transmission of the Schwarz process, only conformal meshes will be used for the differing domains. We have also chosen for this first test to implement the overlapping Schwarz method with a constant time step $dt$ in all domains. Neither of these two specifications are required constraints on the Schwarz method.
@@ 4 Results
[…]
For the purposes of an initial test of the OpInf-Schwarz method, we are most interested if the model matches the standard behavior of ROMs and the Schwarz method. We expect that, as the ROM dimension $r$ and the data allowance increases, the error should approach 0, and that as we increase the size of the Schwarz overlap, the number of Schwarz iterations to converge per timestep will decrease. Lastly, as for any ROM, we expect a significant speed-up in computation time for the price we pay in accuracy.
@@
We will investigate each of these points in a variety of domain decomposition scenarios according to Figure 2. For a global domain, we choose the square $\Omega=[-1,1]\times[-1,1]\in\mathbb{R}^{2}$ on which a monolithic FE simulation will be carried out. The monolithic boundaries will be referred to as $\partial\Omega_{L}$ for the left, $\partial\Omega_{R}$ for the right, $\partial\Omega_{T}$ and $\partial\Omega_{B}$ for the top and bottom of the square respectively. We will divide each square into either two rectangles or four squares, each with an equal Schwarz overlap measured in the number of elements overlapping each neighboring subdomain. For these Schwarz overlaps, explicit communication is only between neighbors which share at least one complete edge – for example, boundary information can be communicated to a left or right neighbor, but not a diagonal neighbor. The Schwarz subdomains will be updated in ascending numerical order.
Aside from the consistent monolithic domain, we will also use the following consistent choices across all tests. We will be using the heat equation as specified in (1) with $f=0$. Our time domain will be $[0,1]$, our time step will be $dt=0.01$ leading to 101 steps, inclusively, from the initial condition at $t=0$ to the final evaluation at $t=1$. The monolithic domain will be split into a 50 by 50 grid, and then triangular Lagrangian elements of degree 1 will be formed by splitting the grid diagonally. These choices lead to 2601 FE nodes on the monolithic mesh, inclusive of the boundary. Within the Schwarz method, our relative (5) and absolute (4) tolerances for convergence will both be fixed at $1\times 10^{-10}$.
@@
$\mathbf{E}_{\ell^{2}}^{\text{avg}}=\frac{1}{\tau-1}\sum_{i=2}^{\tau}\frac{\|\mathbf{u}_{i,\text{MONO}}-\mathbf{u}_{i,\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}_{i,\text{MONO}}\|_{2}},$
Where $\mathbf{u}_{i,\text{MONO}}$ refers to the finite element monolithic simulation carried out over $\Omega$ at time step $i$ and $\mathbf{u}_{i,\text{SCHWARZ}}$ to the solution via the Schwarz method merged from the component subdomains at time step $i$. The norm $\|\cdot\|_{2}$ refers to the $\ell^{2}$ norm, and both $\mathbf{u}_{i,\text{MONO}}$ and $\mathbf{u}_{i,\text{SCHWARZ}}$ are vectors of solution values at each FE node. The error determination is made without including the initial condition, as they are set to be equal.
The projection error of the snapshots onto the POD basis in subdomain $\Omega_{j}$ (dependent on basis dimension $r$) is here defined as,
$\mathbf{E}_{\text{proj}}^{\Omega_{j}}=\frac{\|\mathbf{X}_{j}-\mathbf{\Psi}_{r,j}\mathbf{\Psi}_{r,j}^{T}\mathbf{X}_{j}\|_{F}}{\|\mathbf{X}_{j}\|_{F}},$
where $\mathbf{X}_{j}$ are the snapshots associated with all solved for DoFs in $\Omega_{j}$ over the entire time domain, and $\mathbf{\Psi}_{r,j}$ is the $r$-dimensional basis associated with the solved for DoFs of $\Omega_{j}$. This gives a measure of the approximation quality of the POD basis in each subdomain. When averaged across all subdomains, this will be reported using the symbol, $\mathbf{E}_{\text{proj}}^{\text{avg}}$.
Lastly, we will also use the maximum error in 4.2, defined as,
$\mathbf{E}_{\ell^{2}}^{\text{max}}=\max_{i}\frac{\|\mathbf{u}_{i,\text{MONO}}-\mathbf{u}_{i,\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}_{i,\text{MONO}}\|_{2}}.$
@@
In addition to errors, we will be reporting the following quantities in following tables.
@@
    “Overlap” is the amount by which the $50\times 50$ grid is shared by neighboring subdomains before being broken up into triangular elements.
@@
    “Time” is the average wall clock time taken to run the online model on identical hardware, measured in seconds. The number of runs averaged will be mentioned in each table. This does not include any time spent setting up matrices, generating data, or similar offline tasks.
@@
    “Data” is the number of training steps taken, with a value of 1 only including the initial condition. There are a maximum of 101 steps possible across the time interval $[0,1]$, and any value of “Data” less than 101 will result in prediction outside of training data. “Data” is uniform across all subdomains.
Lastly, computational speed-up will be measured as the wall clock time to run the online portion of the OpInf-Schwarz coupled ROM models relative to that of the fully FE coupled Schwarz method and the monolithic FE solution over the same domain configuration and on the same hardware.

### v3

@@ 2.2.2 Operator Inference
Operator inference is a non-intrusive, data-driven, project- ion-based model order reduction technique which learns low-dimensional operators that can be used to approximate the output of a given FOM, which, in this paper, is the heat equation described in (6). In a classical intrusive Galerkin ROM, the operator $\mathbf{K}_{r}\in\mathbb{R}^{r\times r}$ is the operator which best represents the action of $\mathbf{K}$ in the reduced space of dimension $r$ defined by the basis $\mathbf{\Psi}_{r}$. That is,
$\mathbf{K}_{r}=\mathbf{\Psi}_{r}^{T}\mathbf{K}\mathbf{\Psi}_{r}.$ (11)
This is intrusive because it requires access to the FOM matrix $\mathbf{K}$, and, practically speaking, the code that produced $\mathbf{K}$.
@@
For the problem considered herein, the 2D heat equation, it is straightforward to see that a projection-based ROM preserves the linear algebraic structure of the FOM, (6). With this observation, we now set up the semi-discretized monolithic OpInf ROM problem:
$\dot{\hat{\mathbf{x}}}=\hat{\mathbf{K}}\hat{\mathbf{x}}+\hat{\mathbf{B}}\mathbf{g}.$ (12)
In order to obtain the reduced operators without access to FOM matrices, one solves a regression problem for the reduced matrices $\hat{\mathbf{K}}\in\mathbb{R}^{r\times r}$ and $\hat{\mathbf{B}}\in\mathbb{R}^{r\times m}$. In particular, given $j\leq\tau$ steps of FOM training data $\mathbf{x}_{p}=\mathbf{x}(t_{p})$, $p=1,\dots,j$, minimize,
$\min_{\hat{\mathbf{K}},\hat{\mathbf{B}}}\sum_{p=1}^{j}\|\dot{\hat{\mathbf{x}}}_{p}-\hat{\mathbf{K}}\hat{\mathbf{x}}_{p}-\hat{\mathbf{B}}\mathbf{g}\|_{2}^{2}.$ (13)
Within the training data, we define the ROM representation of the state at time $t_{p}$ to be $\hat{\mathbf{x}}_{p}=\mathbf{\Psi_{r}^{T}}\mathbf{x}_{p}\in\mathbb{R}^{r}$. The time derivative of the ROM state, $\dot{\hat{\mathbf{x}}}_{p}\in\mathbb{R}^{r}$, may not be explicitly available and can instead be estimated from available data using a difference method. The boundary information, $\mathbf{g}\in\mathbb{R}^{m}$, is the same as in the monolithic FOM.
We now briefly contextualize this in a classical Galerkin ROM context. A typical projection-based ROM creates matrices which are in some way representative of those constructed for the FOM, but OpInf does not start with the assumption that these FOM matrices are available. A more natural comparison is between OpInf and the standard Galerkin ROM for the choice of a particular shared basis $\mathbf{\Psi}_{r}$.
It is guaranteed that, under the assumptions that the time-stepping scheme is convergent as $dt\to 0$, the approximation $\dot{\hat{\mathbf{x}}}_{j}$ converges to the true time derivative at time $t_{j}$ as $dt\to 0$, and linear independence of supplied data ($d\geq\tau$, the total number of snapshots), then, for all $0<\epsilon\in\mathbb{R}$ there exists a timestep $dt$ for the FOM and a $r\leq d$ such that,
@@
The same result holds for $\hat{\mathbf{B}}$ [27]. In the case that a technique known as re-projection is used for the fully discretized problem, it can be shown that the learned OpInf ROM system recovers the standard Galerkin projected intrusive ROM [26].
While this overview has focused on the heat equation, we again stress that operator inference is general to low-order polynomial non-linearity, with results previously shown for a polynomial non-linearity of degree three in a one-dimensional (1D) nuclear reactor model in [27], for a 2D single injector combustion model in [21], and for a 3D rotating detonation rocket engine in [9], among others.
2.3 Opinf-Schwarz Method: FOM-ROM Coupling
We are now in a position to state the Opinf-Schwarz method for the heat equation specified in (1). We state the problem formulation for the case of two overlapping subdomains $\Omega_{1}$ and $\Omega_{2}$ with $\Omega_{1}\cup\Omega_{2}=\Omega$ as in Figure 1 where we place a FOM on $\Omega_{1}$, an OpInf ROM on $\Omega_{2}$, and couple the problems via the Schwarz method. The Opinf-Schwarz formulation can be extended easily to arbitrary numbers of subdomains and combinations of model couplings. For simplicity, we assume that we have access to monolithic FOM training data across the entire domain $\Omega$. The dimensional and subdomain notation presented here follows from that defined for the FOM-FOM coupling in Section 2.1.1.
The OpInf problem on $\Omega_{2}$ requires some details. Given $j\leq\tau$ steps of monolithic data $\mathbf{u}(t_{p})\in\mathbb{R}^{N+m},p=1,\dots,j$ extract the unconstrained state representation on $\Omega_{2}$, $\mathbf{x}_{2}(t_{p})\in\mathbb{R}^{N_{2}}$ as well as the the vector $\gamma_{2}(t_{p})$, which is the state information along the Schwarz boundary $\Gamma_{2}$ at time $t_{p}$. The subdomain local boundary on $\Omega_{2}$ is completed by defining the vector $\mathbf{y}_{2}(t_{p})=[\mathbf{g}_{2}\;\gamma_{2}(t_{p})]^{T}\in\mathbb{R}^{m_{2}}$. We carry out POD on the collection of unconstrained snapshots on $\Omega_{2}$, $\mathbf{x}_{2}(t_{p}),p=1,\dots,j$ to obtain the $\Omega_{2}$-local basis $\mathbf{\Psi}_{r,2}\in\mathbb{R}^{N_{2}\times r}$ for a choice of basis dimension $r$. Once the basis is obtained, we solve the regression problem,
$\min_{\hat{\mathbf{K}}_{2},\hat{\mathbf{B}}_{2}}\sum_{p=1}^{j}\|\dot{\hat{\mathbf{x}}}_{2}(t_{p})-\hat{\mathbf{K}}_{2}\hat{\mathbf{x}}_{2}(t_{p})-\hat{\mathbf{B}}_{2}\mathbf{y}_{2}(t_{p})\|_{2}^{2},$ (14)
for the reduced operators $\hat{\mathbf{K}}_{2}\in\mathbb{R}^{r\times r}$ and $\hat{\mathbf{B}}_{2}\in\mathbb{R}^{r\times m_{i}}$ as in Section 2.2.
After setting up the FOM on $\Omega_{1}$ exactly as in Section 2.1.1, the problems may be coupled using the Schwarz method as follows:
$\left\{\begin{array}[]{rcll}\dot{\mathbf{x}}_{1}^{(k+1)}&=&\mathbf{K}_{1}\mathbf{x}_{1}^{(k+1)}+\mathbf{B}_{1}\mathbf{y}_{1}^{(k+1)}\\ \gamma_{1}^{(k+1)}&=&\mathbf{\Psi}_{r,2}\hat{\mathbf{x}}_{2}^{(k)}\big{|}_{\Gamma_{1}},\end{array}\right.$ (15)
and
$\left\{\begin{array}[]{rcll}\dot{\hat{\mathbf{x}}}_{2}^{(k+1)}&=&\hat{\mathbf{K}}_{2}\hat{\mathbf{x}}_{2}^{(k+1)}+\hat{\mathbf{B}}_{2}\mathbf{y}_{2}^{(k+1)}\\ \gamma_{2}^{(k+1)}&=&\mathbf{x}_{1}^{(k+1)}\big{|}_{\Gamma_{2}},\end{array}\right.$ (16)
for Schwarz iteration $k=0,1,2,\dots$ until convergence following (4)-(5). $\mathbf{\Psi}_{r,2}\hat{\mathbf{x}}_{2}^{(k)}$ is the reconstruction of the ROM state $\hat{\mathbf{x}}_{2}^{(k)}$ on the FOM mesh originating from the data source.
The specific formulation described above is valid for conformal meshes only; some interpolation, evaluation or projection scheme would need to be employed for non-conformal meshes depending on the choice of spatial discretization for non-conformal meshes. Creating a ROM-ROM coupling in this setting only requires finding the reduced operators on $\Omega_{1}$ and reconstructing the values found on the Schwarz boundary onto the FOM mesh. We extend this approach to arbitrary numbers of subdomains by traversing the domains in a round-robin fashion and use the most up-to-date Schwarz boundary information available at every Schwarz iteration.
@@ 3 Software Implementations
We now discuss specific choices made in the implementation of the Opinf-Schwarz method for this paper, in particular those which are not essential to the method itself. Our results include Schwarz coupling implementations for the FOM-FOM (Section 2.1.1) FOM-ROM (Section 2.3) and ROM-ROM (Adaptation of Section 2.3) scenarios. In our implementation, the generic FOM now specifically refers to a finite element simulation, and ROM refers to operator inference. All ordinary differential equations (ODEs), such as those presented for the Opinf-Schwarz method in equations (15)-(16), are discretized in time using backward Euler.
@@
We use the open source FE library FEniCSx in Python for all FE simulations. FEniCSx is a FE problem solving environment with the ability to define variational forms close to actual mathematical notation [1], along with various tools [2, 34, 35] to enable the creation of a FE space on which computations can be performed. Boundary conditions are enforced strongly through a lifting method leading to FOM systems similar to (6).
@@
Operator inference is performed with the Python library OpInf, a PyPI Python package available at https://pypi.org/project/opinf/. OpInf is a set of tools to facilitate learning polynomial reduced order models, applied in this paper as described in Section 2.2. The required data is drawn from a monolithic FE simulation on $\Omega$, also performed in FEniCSx. When solving the OpInf regression problem (14), we estimate the required state derivative information $\dot{\hat{\mathbf{x}}}_{2}$ for through a first order backward difference of available state data.
@@
The numerical stability of OpInf models is a significant concern and area of active research. OpInf models are not guaranteed to be stable, but there are a variety of strategies that can help improve stability of reduced models – independently of OpInf, data pre-processing is commonly used in ROMs to create a model for a transformed set of snapshots. In this paper, we use a centering approach where we build a model for snapshot data $\mathbf{x}-\mathbf{\bar{x}}$, the mean-centered snapshots.
@@
Specific to OpInf, we also employ a regression regularization strategy for equation (13). When solving an OpInf regression problem like (13), we add a regularization form which converts the problem into the following form:
$\min_{\hat{\mathbf{K}},\hat{\mathbf{B}}}\sum_{p=1}^{j}\|\dot{\hat{\mathbf{x}}}_{p}-\hat{\mathbf{K}}\hat{\mathbf{x}}_{p}-\hat{\mathbf{B}}\mathbf{g}\|_{2}^{2}+\lambda^{2}\left(\|\hat{\mathbf{K}}\|^{2}_{F}+\|\hat{\mathbf{B}}\|^{2}_{F}\right).$ (17)
By adding this regularization, we penalize the large entries of our ROM matrices $\hat{\mathbf{K}},\hat{\mathbf{B}}$. In our testing, this centering and regularization strategy has been sufficient to produce a stable ROM ODE. For our results in Section 4, all tables and figures have been produced with regularization parameter $\lambda=10^{-2}$. Regularization is a common necessity in operator inference, see for example [21, 31] for related approaches to regularization of OpInf models.
@@
Finally, the Opinf-Schwarz method that couples these problems has been implemented by the authors of this paper in accordance with Section  2.1.1 and Section 2.3. To simplify the boundary transmission of the Schwarz process, only conformal meshes are used for the differing domains. The output of the Schwarz methods (FOM-FOM, ROM-ROM) as presented in Section 2 are unconstrained state vectors on the interiors of $\Omega_{1}$ and $\Omega_{2}$, when what is desired is a single approximation of the monolithic state across $\Omega$. We have made the choice to reconcile the solutions in the following manner.
Taking the FOM-FOM coupling from Section 2.1.1 as an example, the Opinf-Schwarz method produces $\{\mathbf{x}_{1}(t_{p})\}_{p=1}^{\tau}$ and $\{\mathbf{x}_{2}(t_{p})\}_{p=1}^{\tau}$. We define the merged solution vector $\mathbf{u}(t_{p})_{\text{SCHWARZ}}\in\mathbb{R}^{N}$ to be equal to the DoFs of $\mathbf{x}_{i}(t_{p})$ on those entries associated with the interior of $\Omega_{i}\setminus\Omega_{1}\cap\Omega_{2}$ and equal to the known boundary condition $\mathbf{g}$ on entries associated with $\partial\Omega$. For those entries associated with the overlap region $\Omega_{1}\cap\Omega_{2}$, we assign the value of either the entry belonging to $\mathbf{x}_{1}(t_{p})$ or $\mathbf{x}_{2}(t_{p})$ depending on whether the physical node for that entry is closer to the center of $\Omega_{1}$ or $\Omega_{2}$.
@@ 4 Results
For the purposes of an initial test of OpInf-Schwarz, we are most interested if the model matches the standard behavior of ROMs and the Schwarz method. We expect that, as the ROM dimension $r$ and the data allowance increases, the error should approach 0, and that as we increase the size of the Schwarz overlap, the number of Schwarz iterations to converge per timestep will decrease. Lastly, as for any ROM, we expect a significant speed-up in computation time for the price we pay in accuracy.
@@
We investigate each of these points in several different domain decomposition scenarios as shown to Figure 2. For a global domain, we choose the square $\Omega=[-1,1]\times[-1,1]\in\mathbb{R}^{2}$ on which a monolithic FE simulation is carried out. The monolithic boundaries are referred to as $\partial\Omega_{L}$ for the left, $\partial\Omega_{R}$ for the right, and $\partial\Omega_{T}$ and $\partial\Omega_{B}$ for the top and bottom of the domain respectively. We divide the domain into either two rectangles or four squares, each with an equal Schwarz overlap measured in the number of elements overlapping each neighboring subdomain. For these Schwarz overlaps, explicit communication is only between neighbors which share at least one complete edge –hence, in the Four Squares configuration, boundary information can be communicated to a horizontal or vertical neighbor, but not a diagonal neighbor (e.g. $\Omega_{1}$ communicates directly with $\Omega_{2}$ and $\Omega_{3}$ but not $\Omega_{4}$ in Figure 2(c)). The Schwarz subdomains are updated in ascending numerical order.
The specific problem we test the Opinf-Schwarz method on is the heat equation as specified in (1), so we fix the following quantities for the sake of comparison. Our time domain is $[0,1]$, partitioned with a constant time step of $dt=0.01$ leading to 101 steps, inclusively, from the initial condition at $t_{1}=0$ to the final evaluation at $t_{T}=1$. The monolithic domain is be split into a $50\times 50$ grid, and then triangular Lagrangian elements of degree 1 is formed by splitting the grid diagonally. These choices lead to 2601 FE nodes on the monolithic mesh, inclusive of the boundary. Within the Schwarz method, our relative (5) and absolute (4) tolerances for convergence are both be fixed at $10^{-10}$.
@@
$\mathbf{E}_{\ell^{2}}^{\text{avg}}=\frac{1}{\tau-1}\sum_{p=2}^{\tau}\frac{\|\mathbf{u}(t_{p})_{\text{MONO}}-\mathbf{u}(t_{p})_{\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}(t_{p})_{\text{MONO}}\|_{2}},$
As in Section 3, $\mathbf{u}(t_{p})_{\text{SCHWARZ}}$ refers to the solution via the Schwarz method merged from the component subdomains at time $t_{p}$, and $\mathbf{u}(t_{p})_{\text{MONO}}$ refers to the monolithic finite element simulation on $\Omega$ at time $t_{p}$. The error determination excludes the given initial condition, but includes the boundary condition for computational convenience.
The projection error of the snapshots onto the POD basis in subdomain $\Omega_{i}$ (dependent on basis dimension $r$) is here defined as,
$\mathbf{E}_{\text{proj}}^{\Omega_{i}}=\frac{\|\mathbf{X}_{i}-\mathbf{\Psi}_{r,i}\mathbf{\Psi}_{r,i}^{T}\mathbf{X}_{i}\|_{F}}{\|\mathbf{X}_{i}\|_{F}},$
where $\mathbf{X}_{j}$ (chosen to recall the unconstrained state variable $\mathbf{x}$, which the POD basis is built for) are the snapshots associated with all solved for DoFs in $\Omega_{i}$ over the entire time domain, and $\mathbf{\Psi}_{r,i}$ is the $r$-dimensional basis associated with the solved for DoFs of $\Omega_{i}$. This gives a measure of the approximation quality of the POD basis in each subdomain. When averaged across all subdomains, this is reported using the symbol, $\mathbf{E}_{\text{proj}}^{\text{avg}}$.
Lastly, we also use the maximum error in Section 4.2, defined as
$\mathbf{E}_{\ell^{2}}^{\text{max}}=\max_{p}\frac{\|\mathbf{u}(t_{p})_{\text{MONO}}-\mathbf{u}(t_{p})_{\text{SCHWARZ}}\|_{2}}{\|\mathbf{u}(t_{p})_{\text{MONO}}\|_{2}}.$
@@
In addition to errors, we report the following quantities.
@@
    “Overlap” is the amount by which the $50\times 50$ grid is shared by neighboring subdomains before being broken up into triangular elements. For example, in the Four Squares configuration of Figure 2(c), if Overlap = 5, then the subdomain-specific grid for $\Omega_{1}$ shares five rows of the original grid with $\Omega_{3}$ and five columns with $\Omega_{2}$. The overlap with $\Omega_{4}$ is incidental.
@@
    “Time” is the average wall clock time taken to run the online model on identical hardware, measured in seconds. The number of runs averaged are mentioned in each table. This does not include any time spent setting up matrices, generating data, or similar offline tasks.
@@
    “Data” is the number of training steps taken, with a value of 1 only including the initial condition. There are a maximum of 101 steps possible across the time interval $[0,1]$, and any value of “Data” less than 101 results in prediction outside of training data. “Data” is uniform across all subdomains.
Lastly, computational speed-up is measured as the wall clock time to run the online portion of the OpInf-Schwarz coupled ROM models relative to that of the fully FE coupled Schwarz method and the monolithic FE solution over the same domain configuration and on the same hardware.